# Elastic Shape Analysis of Movement Data

### v2

@@ Abstract
Osteoarthritis (OA) is a prevalent degenerative joint disease, with the knee being the most commonly affected joint. Modern studies of knee joint injury and OA often measure biomechanical variables, particularly forces exerted during walking. However, the relationship among gait patterns, clinical profiles, and OA disease remains poorly understood. These biomechanical forces are typically represented as curves over time, but until recently, studies have relied on discrete values (or landmarks) to summarize these curves. This work aims to demonstrate the added value of analyzing full movement curves over conventional discrete summaries. Using data from the Intensive Diet and Exercise for Arthritis (IDEA) study (Messier et al., 2009, 2013), we developed a shape-based representation of variation in the full biomechanical curves. Compared to conventional discrete summaries, our approach yields more powerful predictors of disease severity and relevant clinical traits, as demonstrated by a nested model comparison. Notably, our work is among the first to use movement curves to predict disease measures and to quantitatively evaluate the added value of analyzing full movement curves over conventional discrete summaries.
Keywords: Functional data analysis, Ground reaction forces, Human movement data, Object oriented data analysis, Osteoarthritis, Shape statistics
@@ 1 Introduction
Osteoarthritis (OA) is a prevalent degenerative joint disease characterized by cartilage loss, bone and soft tissue changes, joint pain, and diminished function. In the United States, OA affects at least $19\%$ of adults aged 45 years and older, with the knee being the most frequently affected joint, accounting for more than $80\%$ of the total burden of the disease (Dillon et al., 2006; Jordan et al., 2007; Lawrence et al., 2008; Vos et al., 2012). Knee OA is characterized by both the severity of radiographic damage and clinical symptoms, such as in knee pain and function. Previous research, such as in Zhang and Jordan (2010), indicated that risk factors and disease progression may vary by clinical phenotype. Additionally, important work like Felson (2013) and Guilak (2011) identified biomechanical factors in the etiology and pathogenesis of knee OA.
@@
Biomechanical variables, particularly forces exerted during walking, are often measured in modern studies of joint injury and OA. However, the relationship among gait patterns, clinical profiles, and OA disease remains poorly understood. Biomechanical forces, collected continuously during gait analysis, are typically represented as curves over time. Discrete summaries, such as peak values, are commonly used to analyze patterns in joint loading (forces acting on the knee joint). Such discrete summaries are called landmarks in the shape statistics terminology of Dryden and Mardia (2016). Analyses based on conventional discrete summaries, such as those by Sims et al. (2009) and Astephen et al. (2008), have identified differences between groups (e.g., sex differences) and indicated variations in gait patterns which contribute to knee OA-related outcomes. Recent work by Buck et al. (2024) evaluated the ability of various clinical traits and conventional discrete summaries of gait forces to predict early symptomatic knee OA. While simplifying the statistical methods required for analysis, relying on conventional discrete summaries risks overlooking information encoded in the complete range and patterns of movement data. Research by Muniz et al. (2006), Costello et al. (2021), Bjornsen et al. (2024), Davis et al. (2019), and others has demonstrated the value of analyzing full movement curves. However, these studies are limited and have not been formally compared to analyses based on conventional discrete summaries. For harmonic analyses of full movement curves, see Trentadue and Schmitt (2024) and references therein.
@@
The primary goal of this work is to demonstrate the value added from analyzing full movement curves over conventional discrete summaries. We take an Object Oriented Data Analysis (OODA) approach for this analysis. OODA, described in Marron and Dryden (2021), is a framework for interdisciplinary research that introduces new terminology and methodologies for analyzing complex data. It emphasizes the careful selection of relevant data objects and the utilization of methods intrinsic to the corresponding object spaces. This approach enables the consideration of full movement data curves as complex data objects in high-dimensional space. Furthermore, the richer information within these curves presents numerous choices for both the data object of interest and the appropriate methodology for analysis.
@@
Using data from the Intensive Diet and Exercise for Arthritis (IDEA) study (Messier et al., 2009, 2013), we developed a shape-based representation of variation in the full biomechanical curves. Our approach, compared to conventional discrete summaries, yields more powerful predictors of disease severity and relevant clinical traits, as demonstrated by a nested model comparison. Notably, our work is among the first to use movement curves to predict disease severity and to quantitatively evaluate the added value of utilizing full movement curves as predictors over conventional discrete summaries.
@@ 2 Curves as data objects
For the IDEA study, participants wore laboratory-provided cushioned shoes and walked at their preferred speed on a 22.5m walkway. Kinetic data, including Ground Reaction Force (GRF), were collected using an Advanced Medical Technologies, Inc. model OR-6-5-1 force plate (480 Hz) and filtered using a $4^{\text{th}}$ order low-pass Butterworth filter with a cutoff frequency of 6 Hz. The GRF consists of three components: vertical GRF (vGRF), representing the force exerted downwards; anterior-posterior GRF (apGRF), the propulsive or braking force in the direction of walking; and medial-lateral GRF (mlGRF) in the third orthogonal direction.
When plotted as a function of time, the horizontal axis reflects the percentage of stance (time interval of foot contact with the ground) over a gait cycle with body weight-normalized force values (measured in Newtons) on the vertical axis.
@@
Successful walking trials required the entire foot to be on the force plate during the stance phase, with participants maintaining their preferred walking speed within ±$\pm 3.5\%$. Walking speed was defined as the mean speed over six practice trials, during which participants walked a 10m walkway at a self-selected, stable, and comfortable pace. A photocell system registered the speed, providing immediate visual feedback.
@@
We studied the part of the IDEA data that contained measurements of each GRF component taken at a constant sampling rate over the duration of each step. For each IDEA participant, three trials per limb were collected. For some participants, one or more trials were missing from the data. In those cases, we used as many trials as were given in the data and did not impute. We considered the collection of all trials of both limbs of all participants.
[…]
@@
The top right panel of Figure 1 presents the collection of vGRF curves shown in the first panel, following this processing. Those curves are also color-coded according to their respective walking speeds. Similar results were achieved using the same processing steps for the apGRF and mlGRF curves, and are shown in the bottom left and bottom right panels of Figure 1, respectively. For a more detailed view of GRF curve variations across walking speeds, additional plots showing subsets of the curves grouped by deciles of walking speed are provided in Section 1 of the supplementary material.
@@ 2.1 Conventional discrete summaries
Most relevant studies of GRF data, such as those in Messier et al. (1992), Hunt et al. (2006), Zeni Jr and Higginson (2009), and Wiik et al. (2017), focus on discrete summaries (also called landmarks) rather than analyzing the full curves. These discrete summaries typically correspond to critical subphases of stance during gait, such as the first and second vGRF peaks (heel-strike and toe-off) and positive and negative peaks of the apGRF curve (braking and propulsion). In contrast, mlGRF curves are highly variable and therefore more challenging to interpret and summarize (Costello et al., 2021). Consequently, GRF studies primarily analyze only the vGRF and apGRF directions.
Traditionally, greater attention has been given to the first vGRF peak, as heel-strike is a primary indicator of compressive joint loads and has been consistently linked to knee OA onset and progression. However, recent work by Buck et al. (2024) (and references therein) indicated that the valley (reflecting mid-stance) and the second peak of the vGRF curve are better predictors of knee OA-related symptoms.
In this analysis, we focus on three conventional discrete summaries: (1) the first vGRF peak, (2) the apGRF braking peak, and (3) the apGRF propulsion peak. Throughout the remainder of this paper, we refer to the 3-dimensional vector containing these landmark values as the conventional discrete summary predictors. While other landmark choices could have been analyzed, as discussed in the preceding paragraphs, these three were selected because they are the most commonly studied GRF features in biomechanical research on knee OA.
@@ 3.1 Elastic shape analysis
Elastic warping of the time axis can provide aligned curves that better capture amplitude variation, as studied in Section 2.1 of Marron and Dryden (2021). This is important because poor alignment of curves due to phase variation can impact statistical methodology, potentially obscuring important geometric structure. As noted by Helwig et al. (2011), temporal alignment is a critical consideration in gait analysis, with various approaches being applied in gait studies. That paper highlights that while linear length normalization, which we applied as a pre-processing step, removes differences in stance duration, it does not account for phase variation in the timing of key events (e.g., peaks and valleys). Other methods discussed by Helwig et al. (2011) require extensive manual tuning and may introduce distortions in curve shape. The results discussed in that paper provide evidence for the need for an alignment method that captures phase variation while preserving curve shape and minimizes manual tuning.
@@
Elastic warping involves a transformation of the time axis, which is described by a curve that can be usefully thought of as a stretching and compression of the horizontal axis. The functions are aligned by finding the Karcher mean (Tucker et al., 2013) which produces aligned functions and warping functions and will be defined later on. A warping function: $\gamma(x):[0,1]\rightarrow[0,1]$ is strictly increasing, invertible, and diffeomorphic, in other words the function and its inverse are smooth. The collection of such warping functions serve as phase data objects.
@@
Aligning points across functions is often referred to as registration. Many conventional Functional Data Analysis (FDA) techniques rely on the $\mathbb{L}^{2}$ norm, which simplifies computations into point-wise evaluations. While point-wise computations involve vertical registration, other methods focus on the shape of functions. $L^{2}$-based methods present well-known challenges, as detailed in Marron et al. (2015). Elastic shape analysis, as proposed in Srivastava et al. (2011); Tucker et al. (2013), uses the warp-invariant Fisher-Rao metric to overcome the limitations of conventional $\mathbb{L}^{2}$-based alignment techniques. This framework dates back to such seminal work as Younes (1998) and is the first to enable fully automatic (meaning that no manual tuning is needed) shape-based registration. For each curve, the elastic shape analysis method computes the warp needed to align its peaks to a template mean curve, known as the Karcher mean. The Karcher mean is the curve that lies in the “center” of all the warped curves, and it minimizes the total distance (under the Fisher-Rao metric) between itself and all the aligned curves.
@@
The key idea of this method is to define an equivalence relation between curves. Amplitude (or called shape in Srivastava et al. (2011); Wu et al. (2024)) of the data objects are equivalence classes. Two curves $f_{1}(x)$ and $f_{2}(x)$ are called equivalent, $f_{1}\sim f_{2}$, if there exists a warping function $\gamma$ such that $f_{1}(\gamma(x))=f_{1}\circ\gamma(x)=f_{2}(x).$ Then, the set of all warps of a function $f$, given by
$[f]=\{f\circ\gamma:\gamma\in\Gamma\},$
is the amplitude equivalence class.
@@
The Fisher-Rao metric defines a proper distance on the set of such equivalence classes. In the elastic shape analysis framework, a Square Root Velocity Function (SRVF) representation transforms the Fisher-Rao metric into the standard $\mathbb{L}^{2}$ metric, allowing for a natural framework for the computations required for curve alignment. The Karcher mean equivalence class is defined using this distance, along with the warping functions needed to align individual functions to the Karcher mean template. The SRVF of a function $f\in\mathcal{F}$ is given by $q(t)=\text{sign}(\dot{f}(t))\sqrt{|\dot{f}(t)|}.$ For any time warping of $f$ by $\gamma\in\Gamma$, the SRVF of the warped function is given by $(q\circ\gamma)\sqrt{\dot{\gamma}},$ which we will denote by $q\star\gamma$ for convenience. Then, the warping functions needed to align a collection of curves $f_{1},\ldots,f_{n}$ with corresponding SRVF representations, $q_{1},\ldots,q_{n}$, to the Karcher mean $\mu_{q}^{\ast}$ are computed by solving the optimization problem:
@@
The output of this optimization problem is the Karcher mean $\mu_{q}^{\ast}$ and the set of optimal warping functions $\{\gamma_{i}^{\ast}\}$, and the problem is solved using dynamic programming. To implement this, the time domain $[0,1]$ is discretized into $N$ equally spaced bins, and each warping function $\gamma_{i}$ is approximated by a piecewise linear mapping from $(0,0)$ to $(1,1)$. This discretization corresponds to a path through an $N\times N$ grid, where each step must have a positive slope to ensure invertibility.
In our analysis, we found substantial benefits by penalizing the amount of elasticity in the alignment of the curve. This is achieved by modifying the optimization problem to include a penalty term on the roughness of the $\gamma_{i}$ as follows:
@@
[…]
@@
An interesting alternative to adding a smoothness penalty is a metric learning approach that considers the broader 1-parameter family of elastic metrics, which extends SRVFs as described in Bauer et al. (2024). This family allows for flexible control over warping by varying the transformation in the metric, thereby implicitly enforcing smoothness without requiring an additional penalty term on the $\gamma_{i}$.

### v3

@@ Abstract
Osteoarthritis (OA) is a highly prevalent degenerative joint disease, and the knee is the most commonly affected joint. Biomechanical factors, particularly forces exerted during walking, are often measured in modern studies of knee joint injury and OA, and understanding the relationship among biomechanics, clinical profiles, and OA has high clinical relevance. Biomechanical forces are typically represented as curves over time, but a standard practice in biomechanics research is to summarize these curves by a small number of discrete values (or landmarks). The objective of this work is to demonstrate the added value of analyzing full movement curves over conventional discrete summaries. We developed a shape-based representation of variation in full biomechanical curve data from the Intensive Diet and Exercise for Arthritis (IDEA) study (Messier et al., 2009, 2013), and demonstrated through nested model comparisons that our approach, compared to conventional discrete summaries, yields stronger associations with OA severity and OA-related clinical traits. Notably, our work is among the first to quantitatively evaluate the added value of analyzing full movement curves over conventional discrete summaries.
Keywords: Functional data analysis, Biomechanical data, Nonparametric methods, Shape statistics
@@ 1 Introduction
Osteoarthritis (OA) is a prevalent degenerative joint disease characterized by cartilage loss, bone and soft tissue changes, joint pain, and diminished function. In the United States, OA affects at least $19\%$ of adults aged 45 years and older, with the knee being the most commonly affected joint, accounting for more than $80\%$ of the total burden of the disease (Dillon et al., 2006; Jordan et al., 2007; Lawrence et al., 2008; Vos et al., 2012). Knee OA is characterized by both the severity of radiographically-assessed damage and clinical symptoms, such as knee pain and function. Previous research, such as in Zhang and Jordan (2010), indicated that risk factors and disease progression may vary by clinical phenotype. Additionally, important work like Felson (2013) and Guilak (2011) identified biomechanical factors in the etiology and pathogenesis of knee OA.
@@
Biomechanical variables, particularly forces exerted during walking, are often measured in modern studies of knee joint injury and OA, and understanding the relationship among biomechanics, clinical profiles, and OA has high clinical relevance. Biomechanical forces are typically represented as curves over time, but a standard practice in biomechanics research is to summarize these curves by a small number of discrete values. Such discrete summaries are called landmarks in the shape statistics terminology of Dryden and Mardia (2016). Analyses based on conventional discrete summaries, such as those by Sims et al. (2009) and Astephen et al. (2008), have identified differences between groups (e.g., sex differences) and discovered variations in gait patterns associated with knee OA-related outcomes. Recent work by Buck et al. (2024) evaluated the ability of various clinical traits and conventional discrete summaries of gait forces to predict early symptomatic knee OA. While simplifying the statistical methods required for analysis, relying on conventional discrete summaries risks overlooking information encoded in the complete range and patterns of movement data. Research by Muniz et al. (2006), Davis et al. (2019), Costello et al. (2021), Bjornsen et al. (2024), and others indicate the value of analyzing full movement curves. However, these studies are limited and did not formally compare analyses based on conventional discrete summaries. For harmonic analyses of full movement curves, see Trentadue and Schmitt (2024) and references therein.
@@
The primary goal of this work is to demonstrate the added value of analyzing full movement curves over conventional discrete summaries. Our analysis of full movement curves follows an Object Oriented Data Analysis (OODA) approach. OODA, described in Marron and Dryden (2021), is a framework for analyzing complex data that emphasizes the careful selection of data objects for a given scientific question and the utilization of methods intrinsic to the data object space. This approach facilitates the consideration of full movement data curves as complex data objects in high-dimensional space. Furthermore, the richer information within these curves allows for potentially many different choices of both the data object of interest and the appropriate methodology for analysis.
@@
We developed a shape-based representation of variation in full biomechanical curves using data from the Intensive Diet and Exercise for Arthritis (IDEA) study (Messier et al., 2009, 2013), and demonstrated through easily interpretable nested model comparisons that our approach, compared to conventional discrete summaries, yields stronger associations with OA severity and OA-related clinical traits. Notably, our work is among the first to quantitatively evaluate the added value of analyzing full movement curves over conventional discrete summaries in OA research.
@@ 2 Curves as Data Objects
During gait data collection in the IDEA study, participants wore laboratory-provided cushioned shoes and walked at their preferred speed on a 22.5m walkway. Kinetic data, including Ground Reaction Force (GRF), were collected using an Advanced Medical Technologies, Inc. model OR-6-5-1 force plate (480 Hz) and filtered using a $4^{\text{th}}$ order low-pass Butterworth filter with a cutoff frequency of 6 Hz. The GRF consists of three components: vertical GRF (vGRF), representing the force exerted downwards; anterior-posterior GRF (apGRF), the propulsive or braking force in the direction of walking; and medial-lateral GRF (mlGRF) in the third orthogonal direction. When representing GRF as a function of time, the horizontal axis reflects the percentage of stance (time interval of foot contact with the ground) over a gait cycle, and the vertical axis corresponds to body weight-normalized force values (measured in Newtons).
@@
A walking trial was recorded as a successful observation when the participant’s entire foot maintained contact with the force plate throughout the stance phase, and participants maintained their preferred walking speed within $\pm 3.5\%$. Walking speed was defined as the mean speed at which participants walked a 10m walkway at a self-selected pace over six practice trials. A photocell system registered speed and provided participants with real-time visual feedback for maintaining their walking speed during formal trials.
@@
We studied the part of the IDEA data that contained measurements of each GRF component taken at a constant sampling rate over the duration of each step. For each IDEA participant, three trials per limb were collected. For some participants, one or more trials were missing from the data. In those cases, we used as many trials as were given in the data and did not impute. We considered the collection of all trials of both limbs from all participants.
The complete set of curves for each GRF component contained 2,686 curves from 454 participants. The top left panel of Figure 1 shows the collection of raw data vGRF curves colored by walking speed. There, the rainbow descends from fastest walking speed (red) to slowest (purple). Notice that many of those curves have consecutive starting and trailing zeros, which are outside the stance phase (i.e., after the foot has left contact with the force plate, but force data were still being collected) and hence do not correspond to a meaningful part of the measurements. To account for these spurious measurements, we take the beginning of the force curve as the zero value immediately preceding the first nonzero value, and similarly, the end of the curve as the zero value immediately following the last nonzero value. Using the relevant segment of each curve, we re-scaled the horizontal axis to the unit interval [0,1] in order to establish a common time axis across all curves. Additionally, we applied linear interpolation to the force values, aligning them to an evenly spaced grid. Time normalization to the stance phase $(0–100\%)$ is a standard approach in biomechanics for analyzing GRF and other kinetic variables. The most commonly used method for time normalization is linear length normalization, which is a rescaling of the stance phases to a standard interval (Helwig et al., 2011), as was applied in our analysis. While this approach removes explicit information about stance duration, it retains timing differences in key joint-loading events, such as heel-strike and toe-off, which impact cartilage stress and are therefore relevant to understanding OA. In this work, we chose to focus on a careful analysis of shape, and leave analyzing stance duration as a direction for future work. However, stance duration is strongly correlated with walking speed and distance (Hebenstreit et al., 2015), which were retained as clinical variables in our analysis.
@@
The top right panel of Figure 1 presents the collection of vGRF curves shown in the first panel, following this processing. Those curves are also color-coded according to the participant’s walking speed. Similar results were achieved using the same processing steps for the apGRF and mlGRF curves, and are shown in the bottom left and bottom right panels of Figure 1, respectively. For a more detailed view of GRF curve variations across walking speeds, additional plots showing subsets of the curves grouped by deciles of walking speed are provided in Section 1 of the supplementary material.
@@ 2.1 Conventional discrete summaries
Most relevant studies of GRF data, such as those in Messier et al. (1992), Hunt et al. (2006), Zeni Jr and Higginson (2009), and Wiik et al. (2017), rely on discrete summaries (also called landmarks) rather than analyzing the full curves. These discrete summaries typically correspond to critical subphases of stance during gait, such as the first and second vGRF peaks (heel-strike and toe-off) and positive and negative peaks of the apGRF curve (braking and propulsion). In contrast, mlGRF curves are highly variable and therefore more challenging to interpret and summarize (Costello et al., 2021). Consequently, GRF studies primarily analyze only the vGRF and apGRF directions.
Traditionally, the first vGRF peak has been of greater interest, as heel-strike is a significant factor in joint compressive loads and has been consistently linked to knee OA onset and progression. However, recent work by Buck et al. (2024) (and references therein) indicated that the valley (reflecting mid-stance) and the second peak of the vGRF curve are better predictors of knee OA-related symptoms.
In this analysis, we focus on three conventional discrete summaries: (1) the first vGRF peak, (2) the apGRF braking peak, and (3) the apGRF propulsion peak. Throughout the remainder of this paper, we refer to the 3-dimensional vector containing these landmark values as the conventional discrete summaries. While other landmark choices could have been analyzed, as discussed in the preceding paragraphs, these three were selected because they are the most commonly studied GRF features in biomechanical research on knee OA.
@@ 3.1 Elastic shape analysis
Elastic warping of the time axis can provide aligned curves that better capture amplitude variation, as studied in Section 2.1 of Marron and Dryden (2021). This is important because poor alignment of curves due to phase variation can impact statistical methodology, potentially obscuring important geometric structure. As noted by Helwig et al. (2011), temporal alignment is a critical consideration in gait analysis, with various approaches applied in gait studies. That paper highlights that while linear length normalization, which we applied as a pre-processing step, removes differences in stance duration, it does not account for phase variation in the timing of key events (e.g., peaks and valleys). Other methods discussed by Helwig et al. (2011) require extensive manual tuning and may introduce distortions in curve shape. The results discussed in that paper support the need for an alignment method that captures phase variation while preserving curve shape and minimizes manual tuning.
@@
Elastic warping involves a transformation of the time axis, which is described by a curve that can be usefully thought of as a stretching and compression of the horizontal axis. The functions are aligned by finding the Karcher mean (Tucker et al., 2013) which produces aligned functions and warping functions and will be defined later on. A warping function $\gamma(x):[0,1]\rightarrow[0,1]$ is strictly increasing, invertible, and diffeomorphic, meaning the function and its inverse are smooth. The collection of such warping functions serve as phase data objects.
@@
Aligning points across functions is often referred to as registration. Many conventional Functional Data Analysis (FDA) techniques rely on the $\mathbb{L}^{2}$ norm, which simplifies computations into point-wise evaluations. While point-wise computations involve vertical registration, other methods focus on the shape of functions. $L^{2}$-based methods present well-known challenges, as detailed in Marron et al. (2015). Elastic shape analysis, as proposed in Srivastava et al. (2011); Tucker et al. (2013), uses the warp-invariant Fisher-Rao metric to overcome the limitations of conventional $\mathbb{L}^{2}$-based alignment techniques. This framework dates back to such seminal work as Younes (1998) and is the first to enable fully automatic (meaning that no manual tuning is needed) shape-based registration. For each curve, the elastic shape analysis method computes the warping function needed to align its peaks to a template mean curve, known as the Karcher mean. The Karcher mean is the curve that lies in the “center” of all the warped curves, meaning it minimizes the total distance (under the Fisher-Rao metric) between itself and all the aligned curves.
@@
The key idea of this method is to define an equivalence relation between curves. Two curves $f_{1}(x)$ and $f_{2}(x)$ are called equivalent, $f_{1}\sim f_{2}$, if there exists a warping function $\gamma$ such that $f_{1}(\gamma(x))=f_{1}\circ\gamma(x)=f_{2}(x).$ Then, the set of all warps of a function $f$, given by
$f]=\{f\circ\gamma:\gamma\in\Gamma\},$
is an equivalence class and defines the amplitude ( called shape in Srivastava et al. (2011); Wu et al. (2024)) of $f$.
@@
The Fisher-Rao metric defines a proper distance on the set of such equivalence classes. A natural framework for carrying out the computations required for curve alignment is achieved through a Square Root Velocity Function (SRVF) representation, which transforms the Fisher-Rao metric into the standard $\mathbb{L}^{2}$ metric. The Karcher mean equivalence class is defined using this distance, along with the warping functions needed to align individual functions to the Karcher mean template. The SRVF of a function $f\in\mathcal{F}$ is given by $q(t)=\text{sign}(\dot{f}(t))\sqrt{|\dot{f}(t)|}.$ For any time warping of $f$ by $\gamma\in\Gamma$, the SRVF of the warped function is given by $(q\circ\gamma)\sqrt{\dot{\gamma}},$ which we will denote by $q\star\gamma$ for convenience. Then, the warping functions needed to align a collection of curves $f_{1},\ldots,f_{n}$ with corresponding SRVF representations, $q_{1},\ldots,q_{n}$, to the Karcher mean $\mu_{q}^{\ast}$ are computed by solving the optimization problem:
@@
The output of this optimization problem is the Karcher mean $\mu_{q}^{\ast}$ and the set of optimal warping functions $\{\gamma_{i}^{\ast}\}$. The optimization the problem is solved using dynamic programming: the time domain $[0,1]$ is discretized into $N$ equally spaced bins, and each warping function $\gamma_{i}$ is approximated by a piecewise linear mapping from $(0,0)$ to $(1,1)$. This discretization corresponds to a path through an $N\times N$ grid, where each step is required to have a positive slope to ensure invertibility.
In our analysis, we found substantial benefit in penalizing the amount of elasticity in the alignment of the curve. This is achieved by modifying the optimization problem to include a penalty term on the roughness of the $\gamma_{i}$ as follows:
@@
The penalty $\mathcal{R}(\gamma_{i})$, controlled by the constant $\lambda>0$, imposes a constraint on $\gamma_{i}$. In our approach, we restrict the second derivative of $\gamma_{i}$, so that $\mathcal{R}(\gamma_{i})=\int_{0}^{1}\ddot{\gamma_{i}}(t)\,dt$. This places a restriction on the smoothness in $\gamma_{i}$ and has the effect of keeping $\gamma_{i}$ closer to the identity warp $\gamma_{i}(t)=t$, thus regulating the level of elasticity in the alignment. The case $\lambda=0$ is referred to as fully elastic alignment, while $\lambda=\infty$ is the non-elastic case.
@@
An interesting alternative to adding a smoothness penalty is a metric learning approach that considers the broader 1-parameter family of elastic metrics, which extend SRVFs as described in Bauer et al. (2024). This family allows for flexible control over warping by varying the transformation in the metric, thereby implicitly enforcing smoothness without requiring a penalty term on the $\gamma_{i}$.